OpenStudy (anonymous):

solve sinx+sin2x=0 for [0,2pi)

OpenStudy (anonymous):

What's there to solve? X=0,pi,2pi? Is that what you want?

OpenStudy (anonymous):

all the answers between [0,2pi)

OpenStudy (anonymous):

sin2x=2sinxcosx sinx+sin2x=sinx+2sinxcosx= sinx(1+2cosx)=0

OpenStudy (anonymous):

now we need to solve sinx=0 and 1+2cosx=0

OpenStudy (anonymous):

when sinx=0 on a unit circle we get x=0 and x=pi

OpenStudy (anonymous):

when cosx=-1/2 we get x=2pi/3 and x=4pi/3

OpenStudy (anonymous):

so you get four answers total, hope you get it;)

OpenStudy (anonymous):

thanks